# A Unifying Framework for Parallelizing Sequential Models with Linear Dynamical Systems

### v1

@@ Appendix C Experimental Details and Additional Experiments
We implemented our experiments using the Equinox library (Kidger & Garcia, 2021) in JAX (Bradbury et al., 2018). In our experiments that report wall-clock time, we use a stochastic implementation of quasi-Newton iterations (Zoltowski et al., 2025) that estimates the diagonal using the Hutchinson estimator (Hutchinson, 1989); see Section 3.4 of Zoltowski et al. (2025) for details.
@@
C.1 Experimental Details for Section˜3.1
@@
We use 10 random seeds, where the randomness controls the sequence of the $S_{5}$ group elements. Each seed uses a batch size of 16, i.e. 16 different $S_{5}$ word problems are evaluated for each run. For each seed, we time how long it takes for 5 runs of the fixed point solver (sequential, Picard, quasi-Newton, or Newton) to evaluate, and record this mean wall-clock time. In Figure˜1, we then plot the median of these 10 mean wall-clock times. We run on an H100 with 80GB of onboard memory.
C.2 Experimental Details for Section˜3.2
@@
The experiment depicted in Figure˜2 closely follows the experiments in Section 4.1 of Lim et al. (2024) and Section 6.1 Gonzalez et al. (2024). We use 10 random seeds, where the randomness controls the initialization of the GRUs, the random inputs to the GRUs, and Rademacher variables used in the stochastic implementation of quasi-Newton iterations (Zoltowski et al., 2025). The GRUs were initialized following the standard initialization practice in Equinox. Each seed uses a batch size of 16, i.e. 16 different GRU trajectories are evaluated for each run. For each seed, we time how long it takes for 5 runs of the fixed point solver (sequential, Picard, quasi-Newton, or Newton) to evaluate, and record this mean wall-clock time. In Figure˜2, we then plot the median of these 10 mean wall-clock times. We run on an H100 with 80GB of onboard memory.
C.3 Experimental Details for Section˜3.3
The potential $\phi$ used for the experiment depicted in Figure˜3 is the negative log probability of a mixture of Gaussians. Each Gaussian is $D$-dimensional, and has a random covariance matrix drawn from a Wishart distribution. For each fixed-point method, we ran 10 random seeds, where the randomness controls the randomly chosen covariance matrices for the mixture of Gaussians and the random inputs for Langevin dynamics. Each seed uses a batch size of 16, i.e. 16 different Langevin trajectories are evaluated for each run. We use a step size $\epsilon=$1\text{\times}{10}^{-5}$\$ for the discrete Langevin steps. For each seed, we time how long it takes for 5 runs of the fixed point solver (sequential, Picard, quasi-Newton, or Newton) to evaluate, and record this mean wall-clock time. In Figure˜3, we then plot the median of these 10 mean wall-clock times. We run on an H100 with 80GB of onboard memory. In order to get convergence for Newton iterations for longer sequence lengths, we had to run with increased precision, using the highest option for the jax_default_matmul_precision flag. See Section˜C.4 for further discussion of the importance of numerical precision for implementation of these LDS-based fixed-point methods.
We also include a small additional experiment in Figure˜6: instead of varying the state size dimension, we instead vary $K$, the number of Gaussians that make up the potential $\phi$ determining the Langevin dynamics. We observe qualitatively similar results to the experiment shown in Figure˜3: viz, that Picard and quasi-Newton iterations enjoy similar convergence rates and wall-clock time in settings where the Jacobian is well-approximated by the identity matrix.
@@
Figure 6: In a similar setting as Figure˜3, we instead consider evaluating Langevin dynamics on a potential defined by the negative log probability of a mixture of $K$ anisotropic Gaussians. Here we keep $D=128$ throughout. We observe qualitatively similar behavior to that shown in Figure˜3, where both Picard and quasi-Newton iterations enjoy similar convergence and wall-clock speed.
C.4 Implementation Considerations
@@
Appendix D Extended related work
D.1 Discussion of parallel chord methods
Chapter 7 of Ortega & Rheinboldt (1970) introduces parallel-chord methods for solving nonlinear equations of the form $F(\mathbf{x})=0$ and shows that parallel-chord methods subsume both Newton and Picard iterations. A parallel-chord method $\mathcal{A}$ is an iterative process to solve for $\mathbf{x}$ satisfying $F(\mathbf{x})=0$ using iterations of the form
$\mathbf{x}^{(i+1)}=\mathbf{x}^{(i)}-T^{-1}F(\mathbf{x}^{(i)}).$ (12)
Newton’s method corresponds to
$T=\dfrac{\partial F}{\partial\mathbf{x}}.$
The term “parallel” in this context does not have to do with applying a parallel scan over the sequence length (which we discuss at length in this paper). Instead, “parallel” in “parallel-chord methods” refers to the way in which Newton’s method finds the zero of a function by making a guess for the zero, and then forming a chord that is parallel to the function that the current guess (see Figure˜7). In one-dimension, the linearization is a line (a chord), while in higher-dimensions the linearization is in general a hyperplane. In Newton’s method, the chord/hyperplane is tangent to the function at the current guess, while for other parallel-chord methods the approximate linearization is in general not tangent.
@@ D.1 Discussion of parallel chord methods
While the problem $F(\mathbf{x})=0$ is fully general for nonlinear equations, in the context of this paper on parallelizing Markovian state space models, we consider the special setting where
$F(\mathbf{x})=[x_{1}-f(x_{0}),x_{2}-f(x_{1}),x_{3}-f(x_{2}),\ldots,x_{T}-f(x_{T-1})],$ (13)
and $f$ is the nonlinear transition function, as defined in eq.˜1. Thus, in the context of parallelizing sequences, it follows that
$\dfrac{\partial F}{\partial\mathbf{x}}=\begin{pmatrix}I_{D}&0&0&\ldots&0&0\\ -\frac{\partial f_{2}}{\partial x}(x_{1})&I_{D}&0&\ldots&0&0\\ 0&-\frac{\partial f_{3}}{\partial x}(x_{2})&I_{D}&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&I_{D}&0\\ 0&0&0&\ldots&-\frac{\partial f_{T}}{\partial x}(x_{T-1})&I_{D}\\ \end{pmatrix}$. (14)
When we plug this form of the Jacobian into $T$ in eq.˜12 and simplify, we obtain the linear dynamical system in eq.˜5.
In their treatment of Picard iterations, Ortega & Rheinboldt (1970) consider a more general formulation than that presented in Shih et al. (2023) or in eq.˜9. Instead, similar to the definition presented in Appendix C.2.3 of Gu et al. (2021), Ortega & Rheinboldt (1970) define Picard iterations in the setting where we have removed a linear component of $F$, namely we have written
$F(\mathbf{x})=:T\mathbf{x}-G(\mathbf{x}),$ (15)
for some nonsingular $T$. Note that such redefinition of $F$ in terms of $T$ and $G$ is always possible and not uniquely determined. After making such a redefinition, Ortega & Rheinboldt (1970) define a Picard iteration as an update of the form
$\mathbf{x}^{(i+1)}=T^{-1}G(\mathbf{x}^{(i)}).$ (16)
However, by multiplying both sides of eq.˜15 by $T^{-1}$, it follows that
$T^{-1}G(\mathbf{x}^{(i)})=\mathbf{x}^{(i)}-T^{-1}F(\mathbf{x}^{(i)}),$
showing that the Picard iterations as defined in eq.˜16 fit into the parallel-chord framework set out in eq.˜12. Note that Picard iterations as defined by Shih et al. (2023) or in eq.˜9 of this paper also fit into the framework of eq.˜15: in the context of evaluating discretized ODEs, the residual defined in eq.˜13 becomes
@@
Thus, in the context of eq.˜15, we have that the resulting $G_{t}(\mathbf{x})=\epsilon g_{t-1}(x_{t-1})$, while the resulting $T$ operator is given by
$T=\begin{pmatrix}I_{D}&0&0&\ldots&0&0\\ -I_{D}&I_{D}&0&\ldots&0&0\\ 0&-I_{D}&I_{D}&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&I_{D}&0\\ 0&0&0&\ldots&-I_{D}&I_{D}\\ \end{pmatrix}$.
@@
When we plug this $T$ into eq.˜12 and simplify, we obtain the linear dynamical system in the “Picard” row of Table˜1. In general, the fixed-point methods of the common form given by eq.˜4 all give rise to $T\in\mathbb{R}^{TD\times TD}$ matrices of the form
$T=\begin{pmatrix}I_{D}&0&0&\ldots&0&0\\ -A_{2}&I_{D}&0&\ldots&0&0\\ 0&-A_{3}&I_{D}&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&I_{D}&0\\ 0&0&0&\ldots&-A_{T}&I_{D}\\ \end{pmatrix}$. (17)
@@
Thus, Chapter 7 of Ortega & Rheinboldt (1970) unites Newton and Picard iterations for the general root finding problem $F(\mathbf{x})=0$ under the umbrella of parallel-chord methods, which are iterative updates of the form of eq.˜12. The framework we provide in Table˜1 can be understood as a specialization of parallel-chord methods for the particular problem of sequential evaluation discussed in eq.˜1. Nonetheless, we focus on how in the specific problem of sequential evaluation, which is of great interest in many areas of machine learning, a wide variety of fixed-point methods become iterative application of LDSs, allowing them to be parallelized over the sequence length with an associative scan. This important perspective about parallelizability, which is of great interest in machine learning, is not discussed in Ortega & Rheinboldt (1970) because they are considering a more general problem.
@@
Ortega & Rheinboldt (1970) also discuss in Chapters 7 and 10 of their book how the closeness of the “parallel chord” (in general and in higher dimensions, the “approximating hyperplane”) to the true linearization of the function (Newton’s method) affects the number of iterations needed for the parallel-chord method to converge. In particular, in their Section 7.1, equation 6, Ortega & Rheinboldt (1970) introduce $\sigma(\mathcal{A},F,\mathbf{x}^{*})$, which determines the rate of convergence of $\mathcal{A}$ to the solution $\mathbf{x}^{*}$ of $F(\mathbf{x})$. They define $\sigma$ as
$\sigma(\mathcal{A},F,\mathbf{x}^{*}):=\rho\left(I-T^{-1}\dfrac{\partial F}{\partial\mathbf{x}}(\mathbf{x}^{*})\right),$ (18)
@@
where $\rho(M)$ denotes the spectral radius of a matrix $M$. Of course, if $T=\nicefrac{{\partial F}}{{\partial\mathbf{x}}}(\mathbf{x}^{*})$, then $\sigma=0$. Thus, lower values of $\sigma$ indicates that $T$ is good approximation of the Jacobian matrix $\nicefrac{{\partial F}}{{\partial\mathbf{x}}}$ evaluated at the zero $\mathbf{x}^{*}$ of $F$, while higher values of $\sigma$ indicate that $T$ is a poor approximation for $\nicefrac{{\partial F}}{{\partial\mathbf{x}}}$. Ortega & Rheinboldt (1970) then use $\sigma$ in their Chapter 10 (in particular, their Theorem 10.1.4) to prove linear rates of convergence for parallel-chord methods (where the rate is given by $\sigma$) within a neighborhood of the solution $\mathbf{x}^{*}$.
@@
Thus, a takeaway from Ortega & Rheinboldt (1970) (as paraphrased from Gasilov et al. (1981)) is that the closer $T$ is to $\nicefrac{{\partial F}}{{\partial\mathbf{x}}}$, the fewer iterations are needed for convergence to $\mathbf{x}^{*}$. This takeaway is extremely similar to our guidance, though we specialize to the particular system of equations given by eq.˜3 that results from the goal of rolling out the Markov process given by eq.˜1. Furthermore, instead of focusing on the closeness of $T$ to $\nicefrac{{\partial F}}{{\partial\mathbf{x}}}$ (which in our setting of eq.˜3 are both $TD\times TD$ matrices), we focus on the closeness of a transition matrix $A_{t}$ in an LDS (see the third column of Table˜1) to $\nicefrac{{\partial f}}{{\partial x}}$ (both $D\times D$ matrices).
@@
This difference in focus is important, because in the setting we consider in this paper—using fixed-point iterations of the form eq.˜4 to solve nonlinear equations of the form eq.˜3—Theorem 10.1.4 of Ortega & Rheinboldt (1970) is actually trivial. By “trivial,” we mean that it does not distinguish between the convergence rates of any of the fixed-point iterations we focus on in this paper.
@@
To make this point more precisely, we review [^9] We follow the presentation of Chapter 9 of Ortega & Rheinboldt (1970), in particular Definition 9.2.1. the notion of root-convergence, more commonly known as $R$-convergence.

### v2

@@ Appendix E Experimental Details and Additional Experiments
We implemented our experiments using the Equinox library (Kidger and Garcia, 2021) in JAX (Bradbury et al., 2018). In our experiments that report wall-clock time, we use a stochastic implementation of quasi-Newton iterations (Zoltowski et al., 2025) that estimates the diagonal using the Hutchinson estimator (Hutchinson, 1989); see Section 3.4 of Zoltowski et al. (2025) for details.
@@
E.1 Experimental Details for Section˜C.4
For each scalar LDS with scalar multiplication $\alpha$, we use a state size of $D=2$, i.e. $x_{t}\in\mathbb{R}^{2}$, and a sequence length of $T=100$. We consider 10 random seeds, where the randomness controls the initial state $x_{0}$. We initialize $\mathbf{x}_{1:T}^{(0)}$ at all zeros. We plot the median number of steps to convergence in Figure˜7.
E.2 Experimental Details for Section˜4.1
@@
We use 10 random seeds, where the randomness controls the sequence of the $S_{5}$ group elements. Each seed uses a batch size of 16, i.e. 16 different $S_{5}$ word problems are evaluated for each run. For each seed, we time how long it takes for 5 runs of the fixed point solver (sequential, Picard, quasi-Newton, or Newton) to evaluate, and record this mean wall-clock time. In Figure˜2, we then plot the median of these 10 mean wall-clock times. We run on an H100 with 80GB of onboard memory.
E.3 Experimental Details for Section˜4.2
@@
The experiment depicted in Figure˜3 closely follows the experiments in Section 4.1 of Lim et al. (2024) and Section 6.1 of Gonzalez et al. (2024, Section 6.1). We use 10 random seeds, where the randomness controls the initialization of the GRUs, the random inputs to the GRUs, and Rademacher variables used in the stochastic implementation of quasi-Newton iterations (Zoltowski et al., 2025). The GRUs were initialized following the standard initialization practice in Equinox. Each seed uses a batch size of 16, i.e. 16 different GRU trajectories are evaluated for each run. For each seed, we time how long it takes for 5 runs of the fixed point solver (sequential, Picard, quasi-Newton, or Newton) to evaluate, and record this mean wall-clock time. In Figure˜3, we then plot the median of these 10 mean wall-clock times. We run on an H100 with 80GB of onboard memory.
To demonstrate the different values of the $\mathrm{diff}(\cdot)$ operator for quasi-Newton, Jacobi, and Picard iterations in this setting, we consider the setting $D=8$ and $T=1000$. In Figure˜8, we plot a variety of quantities relevant for the $\gamma$ (cf. eq.˜15) for ten random seeds. We observe that lower values of $\gamma$ (i.e., faster rates of asymptotic linear convergence) coincide with fewer fixed-point iterations needed in Figure˜3.
Figure 8: Understanding the convergence rates in Figure˜3. In the setting of the GRU experiment for $D=8$ and $T=1000$, we plot relevant quantities for understanding the convergence rates of different methods over $10$ random seeds. (Top left.) For all $T=1000$ time steps and over all $10$ random seeds, we plot the spectral norm of the approximate Jacobian for the quasi-Newton iterations we consider in this paper, i.e. $\mathrm{diag}[\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}(x_{t}^{\star})]$. (Top right.) For each of the $10$ random seeds, we plot $\|\widetilde{\mathbf{J}}_{\mathrm{QN}}(\mathbf{x}_{1:T}^{\star})^{-1}\|_{2}$. We observe that they are always larger than one. (Bottom left.) We plot the difference between approximate Jacobians and true dynamics Jacobians over all time steps and seeds for quasi-Newton, Jacobi, and Picard iterations. We observe that this difference for Picard iterations is always larger than one, and so we would intuitively expect Picard iteration to be very slow for parallelizing GRUs. This behavior is precisely what we see in Figure˜3. (Bottom right.) Across the $10$ random seeds, we plot the value of $\gamma$ for Jacobi and quasi-Newton iterations (Picard would be $O(T)$ and so is not shown). Because $\|\widetilde{\mathbf{J}}_{\mathrm{J}}(\mathbf{x}_{1:T}^{\star})^{-1}\|_{2}=1$, the 10 $\gamma_{J}$’s are equivalent to the maximum values from the differences in (bottom left) over the 10 random seeds. However, since (top right) shows that $\|\widetilde{\mathbf{J}}_{\mathrm{QN}}(\mathbf{x}_{1:T}^{\star})^{-1}\|_{2}>1$, we observe that the values of $\gamma_{QN}$ are larger than the values of $\gamma_{\mathrm{J}}$ in (bottom right). In summary, because the values of $\gamma_{J}$ are smaller than the values of $\gamma_{\mathrm{QN}}$, we would intuitively expect Jacobi to converge in fewer fixed-point iterations, which is exactly what we observe in Figure˜3.
We observe that both $\mathrm{diff}(\mathcal{A}_{J})$ and $\mathrm{diff}(\mathcal{A}_{QN})$ are both below one always, which corresponds to their fast rates of convergence demonstrated in Figure˜3. In contrast, $\mathrm{diff}(\mathcal{A}_{P})$ is always greater than one, which corresponds to the slow rates of convergence of Picard iteration in the experiment depicted in Figure˜3.
E.4 Experimental Details for Section˜4.3
@@
Figure 9: In a similar setting as Figure˜4, we instead consider evaluating Langevin dynamics on a potential defined by the negative log probability of a mixture of $K$ anisotropic Gaussians. Here we keep $D=128$ throughout. We observe qualitatively similar behavior to that shown in Figure˜4, where both Picard and quasi-Newton iterations enjoy similar convergence and wall-clock speed.
The potential $\phi$ used for the experiment depicted in Figure˜4 is the negative log probability of a mixture of Gaussians. Each Gaussian is $D$-dimensional, and has a random covariance matrix drawn from a Wishart distribution. For each fixed-point method, we ran 10 random seeds, where the randomness controls the randomly chosen covariance matrices for the mixture of Gaussians and the random inputs for Langevin dynamics. Each seed uses a batch size of 16, i.e. 16 different Langevin trajectories are evaluated for each run. We use a step size $\epsilon=$1\text{\times}{10}^{-5}$\$ for the discrete Langevin steps. For each seed, we time how long it takes for 5 runs of the fixed point solver (sequential, Picard, quasi-Newton, or Newton) to evaluate, and record this mean wall-clock time. In Figure˜4, we then plot the median of these 10 mean wall-clock times. We run on an H100 with 80GB of onboard memory. In order to get convergence for Newton iterations for longer sequence lengths, we had to run with increased precision, using the highest option for the jax_default_matmul_precision flag. See Section˜E.5 for further discussion of the importance of numerical precision for implementation of these LDS-based fixed-point methods.
We also include a small additional experiment in Figure˜9: instead of varying the state size dimension, we instead vary $K$, the number of Gaussians that make up the potential $\phi$ determining the Langevin dynamics. We observe qualitatively similar results to the experiment shown in Figure˜4: viz, that Picard and quasi-Newton iterations enjoy similar convergence rates and wall-clock time in settings where the Jacobian is well-approximated by the identity matrix.
Finally, we include an experiment designed to show how the density of the Jacobian favors Newton iterations over Picard iterations. In the experiment depicted in Figure˜10, we again parallelize discretized Langevin diffusion two-well potentials. However, this time we use three different discretization step-sizes $\epsilon$: we use $1\text{\times}{10}^{-5}$,$1\text{\times}{10}^{-4}$$, and $1\text{\times}{10}^{-3}$. We showed that for step size $\epsilon=$1\text{\times}{10}^{-5}$\$ (the smallest considered, Jacobian most diagonal) that Newton and Picard both converged in a small number of fixed-point iterations, with Picard therefore running faster overall. In contrast, for $\epsilon=$1\text{\times}{10}^{-3}$\$ (the largest considered, Jacobian closer to dense), Newton continued to converge in a small number of iterations whereas Picard required considerably more iterations to converge. In this setting, Newton remained faster than sequential evaluation on wallclock time, whereas Picard did not.
Figure 10: Varying the step size $\epsilon$. In this variant of discretized Langevin diffusion on a two-well potential, we vary the step size $\epsilon$. For the smallest step size $\epsilon=$1\text{\times}{10}^{-5}$\$ the Jacobian is most diagonal dominant of these three settings, and Picard iterations are faster on a wallclock time than Newton iterations. For the largest step size $\epsilon=$1\text{\times}{10}^{-3}$\$ the Jacobian is least diagonal dominant of these three settings, and Picard iterations are slower on a wallclock time than Newton iterations. We do not show Jacobi iterations because they already were demonstrated to struggle in discretizations of Langevin diffusions in Figure˜4.
E.5 Implementation Considerations
@@
Appendix F Extended related work
F.1 Discussion of parallel chord methods
Ortega and Rheinboldt (1970) discuss at length iterative methods for solving arbitrary systems of nonlinear equations $\mathbf{F}(\mathbf{x})=0$ using iterations of the form
$\mathbf{x}^{(i+1)}=\mathbf{x}^{(i)}-\widetilde{\mathbf{J}}(\mathbf{x}^{(i)})^{-1}\mathbf{F}(\mathbf{x}^{(i)})$ (21)
for some matrix $\widetilde{\mathbf{J}}(\mathbf{x}^{(i)})$. In general, $\widetilde{\mathbf{J}}$ can be a function of the current iterate $\mathbf{x}^{(i)}$ or a fixed and constant matrix. Newton’s method corresponds to
$\widetilde{\mathbf{J}}(\mathbf{x}^{(i)})=\dfrac{\partial\mathbf{F}}{\partial\mathbf{x}}(\mathbf{x}^{(i)}).$
When $\widetilde{\mathbf{J}}$ is fixed and constant, Ortega and Rheinboldt (1970) describe the resulting family of fixed-point iterations as parallel-chord methods. However, we will use this term for all iterative methods with updates of the form in eq.˜21, which includes both Newton and Picard iterations.
The term “parallel” in this context does not have to do with applying a parallel scan over the sequence length (which we discuss at length in this paper). Instead, “parallel” in “parallel-chord methods” refers to the way in which Newton’s method finds the zero of a function by making a guess for the zero, and then forming a chord that is parallel to the function at the current guess (see Figure˜11). In one-dimension, the linearization is a line (a chord), while in higher-dimensions the linearization is in general a hyperplane. In Newton’s method, the chord/hyperplane is tangent to the function at the current guess, while for other parallel-chord methods the approximate linearization is in general not tangent.
@@ F.1 Discussion of parallel chord methods
The equation $\mathbf{F}(\mathbf{x})=\mathbf{0}$ is a fully general way to represent a system of nonlinear equations. However, in this paper, we focus on parallelizing Markovian state space models. Thus, we consider the special setting where
$\mathbf{F}(\mathbf{x})=\mathbf{r}(\mathbf{x}_{1:T}):=[x_{1}-f(x_{0}),x_{2}-f(x_{1}),x_{3}-f(x_{2}),\ldots,x_{T}-f(x_{T-1})],$ (22)
where $f$ is the nonlinear transition function, as defined in eq.˜1, and $\mathbf{r}$ is the residual function defined in eq.˜10. Thus, in the context of parallelizing sequences, it follows that
$\dfrac{\partial\mathbf{F}}{\partial\mathbf{x}}(\mathbf{x})=\begin{pmatrix}I_{D}&0&0&\ldots&0&0\\ -\frac{\partial f_{2}}{\partial x}(x_{1})&I_{D}&0&\ldots&0&0\\ 0&-\frac{\partial f_{3}}{\partial x}(x_{2})&I_{D}&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&I_{D}&0\\ 0&0&0&\ldots&-\frac{\partial f_{T}}{\partial x}(x_{T-1})&I_{D}\\ \end{pmatrix}$. (23)
When we plug this form of the Jacobian into $\widetilde{\mathbf{J}}$ in eq.˜21 and simplify, we obtain the linear dynamical system in eq.˜5, i.e. Newton’s method.
In their treatment of Picard iterations, Ortega and Rheinboldt (1970) consider a more general formulation than that presented in Shih et al. (2023) or in eq.˜8. Instead, similar to the definition presented in Appendix C.2.3 of Gu et al. (2021), Ortega and Rheinboldt (1970) define Picard iterations in the setting where we have removed a linear component of $\mathbf{F}$, namely we have written
$\mathbf{F}(\mathbf{x})=:\widetilde{\mathbf{J}}\mathbf{x}-\mathbf{G}(\mathbf{x}),$ (24)
for some constant, nonsingular matrix $\widetilde{\mathbf{J}}$ and nonlinear function $\mathbf{G}(\cdot)$. Note that such redefinition of $\mathbf{F}(\cdot)$ in terms of $\widetilde{\mathbf{J}}$ and $\mathbf{G}(\cdot)$ is always possible and not uniquely determined. After making such a redefinition, Ortega and Rheinboldt (1970) define a Picard iteration as an update of the form
$\mathbf{x}^{(i+1)}=\widetilde{\mathbf{J}}^{-1}\mathbf{G}(\mathbf{x}^{(i)}).$ (25)
However, by multiplying both sides of eq.˜24 by $\widetilde{\mathbf{J}}^{-1}$, it follows that
$\widetilde{\mathbf{J}}^{-1}\mathbf{G}(\mathbf{x}^{(i)})=\mathbf{x}^{(i)}-\widetilde{\mathbf{J}}^{-1}\mathbf{F}(\mathbf{x}^{(i)}),$
showing that the Picard iterations as defined in eq.˜25 fit into the parallel-chord framework set out in eq.˜21. Note that Picard iterations as defined by Shih et al. (2023) or in eq.˜8 of this paper also fit into the framework of eq.˜24: in the context of evaluating discretized ODEs, the residual defined in eq.˜22 becomes
@@
Thus, in the context of eq.˜24, we have that the resulting $G_{t}(\mathbf{x})=\epsilon g_{t-1}(x_{t-1})$, while the resulting $\widetilde{\mathbf{J}}$ operator is given by eq.˜20.
@@
When we plug this $\widetilde{\mathbf{J}}$ into eq.˜21 and simplify, we obtain the linear dynamical system in the “Picard” row of Table˜1. In general, the fixed-point methods of the common form given by eq.˜4 all give rise to $\widetilde{\mathbf{J}}\in\mathbb{R}^{TD\times TD}$ matrices of the form shown in eq.˜19.
@@
Thus, Ortega and Rheinboldt (1970) unites Newton and Picard iterations for the general root finding problem $F(\mathbf{x})=0$ under the umbrella of parallel-chord methods, which are iterative updates of the form of eq.˜21. The framework we provide in Table˜1 can be understood as a specialization of parallel-chord methods for the particular problem of sequential evaluation discussed in eq.˜1. Nonetheless, we focus on how in the specific problem of sequential evaluation, which is of great interest in many areas of machine learning, a wide variety of fixed-point methods become iterative application of LDSs, allowing them to be parallelized over the sequence length with an associative scan. This important perspective about parallelizability, which is of great interest in machine learning, is not discussed in Ortega and Rheinboldt (1970) because they are considering a more general problem.
@@
Ortega and Rheinboldt (1970) also discuss in their Chapters 7 and 10 how the closeness of the “parallel chord” (in general and in higher dimensions, the “approximating hyperplane”) to the true linearization of the function (Newton’s method) affects the number of iterations needed for the parallel-chord method to converge. This analysis is directly analogous to our study of the effect of $\left\|\widetilde{\mathbf{J}}(\mathbf{x}_{1:T})-\mathbf{J}(\mathbf{x}_{1:T})\right\|_{2}$ on the rate of convergence of fixed-point methods, see Lemma˜1 In particular, in Chapter 10 of Ortega and Rheinboldt (1970), they consider the rates of convergence of fixed-point methods with updates taking the form of
$\mathbf{x}^{(i+1)}=\mathbf{U}(\mathbf{x}^{(i)}),$ (26)
for some function $\mathbf{U}(\cdot)$. Ortega and Rheinboldt (1970) use the name one-step stationary methods such fixed-point methods with updates with the form eq.˜26.
For parallel-chord methods of the form given in eq.˜21, it follows that
$\mathbf{U}(\mathbf{x}^{(i)})=\mathbf{x}^{(i)}-\widetilde{\mathbf{J}}(\mathbf{x}^{(i)})^{-1}\mathbf{F}(\mathbf{x}^{(i)}).$ (27)
In particular, in their Chapters 7 and 10, Ortega and Rheinboldt (1970) introduce and study $\sigma(\mathbf{U},\mathbf{F},\mathbf{x}^{\star})$, which determines the rate of convergence of iterative methods with updates of the form given by eq.˜26 to the solution $\mathbf{x}^{\star}$ of $\mathbf{F}(\mathbf{x})=\mathbf{0}$. They define $\sigma$ as
$\sigma(\mathbf{U},\mathbf{F},\mathbf{x}^{\star}):=\rho\left(\dfrac{\partial\mathbf{U}}{\partial\mathbf{x}}(\mathbf{x}^{\star})\right),$ (28)
where $\rho(M)$ denotes the spectral radius of a matrix $M$.
In the context of parallel-chord methods where $\mathbf{U}(\cdot)$ is given by eq.˜27, it follows that
$\dfrac{\partial\mathbf{U}}{\partial\mathbf{x}}(\mathbf{x}^{\star})=\mathbf{I}-\widetilde{\mathbf{J}}(\mathbf{x}^{\star})^{-1}\dfrac{\partial\mathbf{F}}{\partial\mathbf{x}}(\mathbf{x}^{\star}),$
@@
because $\mathbf{F}(\mathbf{x}^{\star})=0$. Thus it follows that if $\widetilde{\mathbf{J}}=\nicefrac{{\partial\mathbf{F}}}{{\partial\mathbf{x}}}(\mathbf{x}^{\star})$, then $\sigma=0$. Thus, lower values of $\sigma$ indicate that $\widetilde{\mathbf{J}}$ is a good approximation of the Jacobian matrix $\nicefrac{{\partial\mathbf{F}}}{{\partial\mathbf{x}}}$ evaluated at the zero $\mathbf{x}^{\star}$ of $\mathbf{F}$, while higher values of $\sigma$ indicate that $\widetilde{\mathbf{J}}$ is a poor approximation for $\nicefrac{{\partial\mathbf{F}}}{{\partial\mathbf{x}}}$. Ortega and Rheinboldt (1970) then use $\sigma$ in their Chapter 10 (in particular, their Theorem 10.1.4) to prove linear rates of convergence101010where the rate is given by $\sigma$ for one-step stationary methods within a neighborhood of the solution $\mathbf{x}^{\star}$.
@@
Thus, a takeaway from Ortega and Rheinboldt (1970) (as paraphrased from Gasilov et al. (1981)) is that the closer $\widetilde{\mathbf{J}}$ is to $\nicefrac{{\partial\mathbf{F}}}{{\partial\mathbf{x}}}$, the fewer iterations are needed for convergence to $\mathbf{x}^{\star}$. This takeaway is extremely similar to our guidance, though we specialize to the particular system of equations given by eq.˜3 that results from the goal of rolling out the Markov process given by eq.˜1.
@@
However, in the setting we consider in this paper—using fixed-point iterations of the form eq.˜4 to solve nonlinear equations of the form eq.˜3—Theorem 10.1.4 of Ortega and Rheinboldt (1970) is actually trivial. By “trivial,” we mean that it does not distinguish between the convergence rates of any of the fixed-point iterations we focus on in this paper.
@@
To make this point more precisely, we review [^11] We follow the presentation of Chapter 9 of Ortega and Rheinboldt (1970), in particular Definition 9.2.1. the notion of root-convergence, more commonly known as $R$-convergence.